OpenStudy (ttp):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 13 breakfasts and used 21 slices of bread. Use a model to determine how many egg sandwiches they sold. 6 7 8 9

OpenStudy (ttp):

I think it's C.8?

OpenStudy (anonymous):

S + E = #scrambled + #egg = 13 1 S + 2 E = # slices of toast used =21 subtract first from second E = 8 Thus S = 5 Check S + 2E =21 5 + 2(8) = 21 right

OpenStudy (zale101):

try using this two variables 'p' is the number of sold scrambled egg plates 's' is the number of egg sandwiches sold so p + ys = 13 breakfasts one slice of p + two sloces of toast s = 21 slices 1p+2s=21 p = 13 - s then use it this equation to substitute it into the second one 1(13 - s) + 2s = 21 now solve for s